# Adapt-Pruner: Adaptive Structural Pruning for Efficient Small Language Model Training

### v2

@@ 4.4 Ablation Study: Optimal Interleaving Frequency
All experiments are conducted on MobileLLM-350M, with 1B tokens sampled from Slimpajama (Soboleva et al., 2023) and 1/6 random samples in the aforementioned dataset (Table 2). Two types of benchmarks are adopted, including the commonsense benchmarks in Section 4.1 and MMLU benchmark in Section 4.2. Pruning ratios per training, ranging from , are searched to decide the optimal value, which corresponds to the number of interleaves separately.
@@ Appendix A Experimental Details
We extend the LLM-Pruner framework [Ma et al., 2023, Fang et al., 2023] as our baseline, which incorporates modules for computing similarity scores and adaptive pruning using PyTorch [Paszke et al., 2019]. Our experiments utilize both NVIDIA GH200 and H100 GPUs, where a single GPU (either GH200 or H100) is utilized for pruning and evaluation tasks while 4 additional GPUs are employed in parallel for post-training. The metrics and the number of shots used for each benchmark are available in Table 6.
@@
For Adapt-Accel and Adapt-LLMs, we keep the same choice of sparsity amplitude in Equation 6. The global batch size is set to 128, with maximal learning rate of , and minimal learning rate of . WSD scheduler [Hu et al., 2024] is applied, where the combination of learning rate in all interleaved training iterations forms a WSD scheduler, with 5% warmup steps and 10% linear decay steps at the end. For the ShearedLLaMA [Xia et al., 2024] comparison, we keep these hyperparameters while allocating 10% of tokens for pruning optimization and 90% for post-training. Similarly, for Minitron comparison [Sreenivas et al., 2024], we keep the same settings and fine-tune the teacher model for one epoch on the dataset.
@@ B.1 Adapt-Pruner as Strong Pruners
We evaluate various structured pruning methods on three LLaMA-series models: LLaMA-3.1-8B, LLaMA-3.2-3B, and LLaMA-3.2-1B [AI@Meta, 2024]. For each model, we test sparsity levels of 20%, 40%, and 60% over three trials to assess our method’s effectiveness across different model sizes.

### v3

@@ 4.4 Ablation Study: Optimal Interleaving Frequency
All experiments are conducted on MobileLLM-350M, with 1B tokens sampled from Slimpajama (slimpajama) and 1/6 random samples in the aforementioned dataset (Table 2). Two types of benchmarks are adopted, including the commonsense benchmarks in Section 4.1 and MMLU benchmark in Section 4.2. Pruning ratios per training, ranging from , are searched to decide the optimal value, which corresponds to the number of interleaves separately.
@@ Appendix A Experimental Details
We extend the LLM-Pruner framework [llm-pruner, depgraph] as our baseline, which incorporates modules for computing similarity scores and adaptive pruning using PyTorch [paszke2019pytorch]. Our experiments utilize both NVIDIA GH200 and H100 GPUs, where a single GPU (either GH200 or H100) is utilized for pruning and evaluation tasks while 4 additional GPUs are employed in parallel for post-training. The metrics and the number of shots used for each benchmark are available in Table 6.
@@
For Adapt-Accel and Adapt-LLMs, we keep the same choice of sparsity amplitude in Equation 6. The global batch size is set to 128, with maximal learning rate of , and minimal learning rate of . WSD scheduler [wsd] is applied, where the combination of learning rate in all interleaved training iterations forms a WSD scheduler, with 5% warmup steps and 10% linear decay steps at the end. For the ShearedLLaMA [xia2024shearedllamaacceleratinglanguage] comparison, we keep these hyperparameters while allocating 10% of tokens for pruning optimization and 90% for post-training. Similarly, for Minitron comparison [sreenivas2024llmpruningdistillationpractice], we keep the same settings and fine-tune the teacher model for one epoch on the dataset.
@@ B.1 Adapt-Pruner as Strong Pruners
We evaluate various structured pruning methods on three LLaMA-series models: LLaMA-3.1-8B, LLaMA-3.2-3B, and LLaMA-3.2-1B [llama3modelcard]. For each model, we test sparsity levels of 20%, 40%, and 60% over three trials to assess our method’s effectiveness across different model sizes.